The molar volume (cm3 /mol) of a binary liquid mixture at a specific T and P is given by: V = 200x1 + 300x2 + x1 x2 (20x1 + 10x2) (a) Find expressions for the partial molar volumes of species 1 and species 2 at specific T and P (b) Show that the expressions for partial molar volumes follow the summability relation (c) Show that these expressions satisfy the Gibbs/Duhem equation at constant T and P (d) Find the values for the followings – ?1; ?2 ; ̅?1 ∞ ; ?̅ 2 ∞
Answer #1

. Sel V = 200k + 300x2 + x1x2 (2014 +loxo) V=Joox +300-30014 + (xu-ki) (son Ho-loxe) V= $ 300- looxy + (04-x)( 10.+10x) V = 3M = Luni (Summobilly Ear) V=xVitrato V= x( 910-30xi +8054?) +(1-1) ( 300+70x19) v=g1ox-30x3+ gor +3.00 + hone-300x1-soxa V =

using guven data and relevant formulae we can find the required variables

for any doubts write in comments